# Weyl metrizability of 3-dimensional projective structures and CR submanifolds

### v2

@@ 2.4. Twistor CR manifolds and their Petrov type
Define $\mathcal{T}^{[g]}=\mathbb{P}_{+}T^{*}M,$ where $\mathbb{P}_{+}T^{*}M$ denotes the bundle of positive rays in $T^{*}M$ with respect to $[g].$ Therefore, in the Riemannian signature $\mathbb{P}_{+}T^{*}M$ is the two-fold cover of $\mathbb{P}T^{*}M$, while in the Lorentzian signature it denotes the rays of future time-like 1-forms. Thus, the twistor bundle $\widetilde{\mu}\colon\mathcal{T}^{[g]}\to M$ is a 5-manifold whose fibers are the 2-sphere, $\mathbb{S}^{2},$ if $[g]$ is Riemannian or the 2-disk, $\mathbb{D}^{2},$ if $[g]$ is Lorentzian. To view $\hat{\mu}\colon\mathcal{C}\to\mathcal{T}^{[g]}$ as an associated bundle of $\mu\colon\mathcal{C}\to M,$ recall that $(\mathcal{C}\to M,\psi)$ has type $(\mathrm{SO}^{o}(p+1,q+1),Q_{1}),$ where $(p,q)$ is either (3,0) or (2,1), $Q_{1}=Q_{0}\ltimes\mathbb{R}^{3},$ $Q_{0}=\mathrm{CO}^{o}(p,q)\cong R^{+}\times\mathrm{SO}^{o}(p,q)$ is the structure group, $\mathrm{SO}^{o}(2,1)$ is the identity component of $\mathrm{SO}(2,1)$ and $\mathrm{SO}^{o}(3,\mathbb{R})=\mathrm{SO}(3,\mathbb{R}).$ The stabilizer of a positively oriented (time-like) plane defines a subgroup $S\subset Q_{1}.$ Thus, one obtains that $\hat{\mu}\colon\mathcal{C}\to\mathcal{T}^{[g]}$ is a principal $S$-bundle i.e.
(2.14) $\mathcal{T}^{[g]}\cong\mathcal{C}/S:=\mathcal{C}\times_{Q_{1}}(Q_{1}/S),\quad\text{where}\quad S=\mathrm{CO}(2,\mathbb{R})\ltimes\mathbb{R}^{3}.$
@@
(2.15) $\mathcal{T}^{[g]}_{x}=Q_{1}/S=\mathrm{SO}^{o}(p,q)/\mathrm{SO}(2,\mathbb{R})\cong\left\{\begin{aligned} &\mathbb{S}^{2}\quad(p,q)=(3,0),\\ &\mathbb{D}^{2}\quad(p,q)=(2,1).\\ \end{aligned}\right.$
@@
In Lorentzian signature, the action of $Q_{1}$ on each oriented projectivized cotangent space has five orbits given by future and past rays of time-like 1-forms, rays of space-like 1-forms, and rays of future and past null 1-forms. One can define a twistor bundle using space-like planes in each tangent space equipped with their canonically induced almost para-complex structure. The construction goes through as before with the only difference being the fibers which are $\mathrm{SO}(2,1)/\mathrm{SO}(1,1)\cong\mathbb{R}\times\mathbb{R}\mathbb{P}^{1}.$
As was mentioned above, $\mathcal{T}^{[g]}$ was originally defined as a connected component of the set of complexified null 2-planes without any real null 2-planes, defined in (2.5), i.e. $\mathcal{N}^{\mathbb{C}}$ in the Riemannian signature and a connected component of $\mathcal{N}^{\mathbb{C}}\backslash\mathcal{N}^{\mathbb{R}}$ in the Lorentzian signature. The two definitions are equivalent since there is a bijection between (time-like) planes equipped with their canonically induced almost complex structure and projective classes of complex null 1-forms $[\omega_{z}]:=[z_{0}\omega^{0}+z_{1}\omega^{1}+z_{2}\omega^{2}]$ in (2.5). To see this, note that given a positive (time-like) plane $p\in\mathrm{Gr}^{+}(2,T_{x}M)$, one can find a coframe $(\omega^{0},\omega^{1},\omega^{2})$ that satisfies (2.2) and $p=\mathrm{Ker}\{\omega^{0}\}.$ The induced $\mathrm{SO}(2,\mathbb{R})$ action on $(\omega^{1},\omega^{2})^{t}$ defines an almost complex structure whose dual $\mathrm{i}$-eigenspace is $\omega^{1}+\mathrm{i}\omega^{2},$ which is a null complex 1-form. Conversely, given a projective class of a complex null 1-form, $[\omega_{z}]:=[z_{0}\omega^{0}+z_{1}\omega^{1}+z_{2}\omega^{2}],$ $z_{0}^{2}+\varepsilon z_{1}^{2}+\varepsilon z_{2}^{2}=0,$ one can find $[A^{i}_{j}]\in\mathrm{SO}^{o}(p,q)$ such that $[\omega_{z}]=[\tilde{\omega}^{1}+\mathrm{i}\tilde{\omega}^{2}]$ where $\tilde{\omega}^{i}=A^{i}_{j}\omega^{j}.$ More explicitly, examining the case $(p,q)=(3,0)$ over the affine chart $z_{2}=\mathrm{i},$ one has $[\omega_{z}]=[z_{0}\omega^{0}+z_{1}\omega^{1}+\mathrm{i}\omega^{2}]$ and $z_{0}^{2}+z_{1}^{2}=1.$ Using the 2-1 isogeny $\mathrm{SU}(2)\to\mathrm{SO}(3,\mathbb{R}),$ the coframe transformation $\omega^{i}\to\tilde{\omega}^{i}$ is
@@
where $|a|^{2}+|b|^{2}=1.$ Finding $a,b\in\mathbb{C}$ such that $[\tilde{\omega}^{1}+\mathrm{i}\tilde{\omega}^{2}]=[z_{0}\omega^{0}+z_{1}\omega^{1}+\mathrm{i}\omega^{2}]$ results in a set of polynomial relations that can be easily solved for any $z_{0},z_{1}\in\mathbb{C}$ satisfying $z_{0}^{2}+z_{1}^{2}=1.$ This shows that $\omega_{z}$ gives the dual $\mathrm{i}$-eigenspace for the almost complex structure on $\mathrm{Ker}\{\tilde{\omega}^{0}\}.$
@@
To show that $\mathcal{T}^{[g]}$ carries a CR structure, we follow [SY89], which is different from [LeB84] and is based on exploiting the embedding of Lie algebras $\mathfrak{so}(p+1,q+1)\hookrightarrow\mathfrak{su}(2,2),$ where $p+q=3,$ given explicitly in (2.16). Firstly, using the description $\mathcal{T}^{[g]}=\mathcal{C}/S$ in (2.14) together with the Cartan connection (2.1), one obtains that the 1-forms $(\omega^{0},\omega^{1},\omega^{2},\theta^{1},\theta^{2})$ are semi-basic with respect to the fibration $\mathcal{C}\to\mathcal{T}^{[g]}.$ Furthermore, by the structure equations (2.3), one knows
@@
where $\omega=\omega^{1}+\mathrm{i}\omega^{2}$ and $\theta=\theta^{1}+\mathrm{i}\theta^{2}.$ The conformal class of Hermitian inner products $[h]\subset\mathrm{Sym}^{2}(\mathscr{C}^{*})$ is given by $h=\tfrac{\mathrm{i}}{2}(\omega\overline{\theta}-\overline{\omega}\theta).$ Using an adapted coframe $(\omega^{0},\omega,\theta,\overline{\omega},\overline{\theta}),$ one can find the Cartan connection for the induced nondegenerate CR structure on $\mathcal{T}^{[g]}.$ Namely, the Lie algebra embedding $\mathfrak{so}(p+1,q+1)\hookrightarrow\mathfrak{su}(2,2)$ results in the injection $\iota\colon S\to P,$ where $P\in\mathrm{SU}(2,2)$ is the parabolic subgroup in Theorem 2.5 and $S$ is given in (2.14). Defining $\mathcal{Q}=\mathcal{C}\times_{S}P,$ one obtains a natural inclusion $\iota\colon\mathcal{C}\to\mathcal{Q}$ between the Cartan geometries $(\mathcal{C}\to M,\psi)$ and $(\mathcal{Q}\to\mathcal{T}^{[g]},\gamma)$ via which
@@
are defined in terms of the entries of the Cartan connection (2.1) for the 3-dimensional conformal structure $[g]$. Using $\iota^{*}\gamma$ and the structure equations (2.3) and (2.8), the coefficients $C_{0},\cdots,C_{4}$ of the harmonic binary quartic (2.9) on $\mathcal{T}^{[g]},$ pulled-back via $\iota,$ are found to be
@@
The Cotton-York entry $Y_{012}$ is identically zero everywhere on $\mathcal{C}$ if and only if the conformal structure is flat. As a result, the CR structure on $\mathcal{T}^{[g]}$ is flat if and only if the conformal structure $[g]$ is flat. Otherwise, from (2.18), it follows that the binary quartic (2.9) has a repeated root of multiplicity four. Using the parametrization of $\alpha$-planes (2.11) and the quartic (2.12), it follows that the repeated root corresponds to the $\alpha$-plane for which $[a_{1}:a_{2}]=[0:1],$ i.e. the contact planes $\mathrm{Ker}\{\omega^{0},\omega^{1},\omega^{2}\},$ whose integral surfaces are the fibers of $\mathcal{T}^{[g]}\to M.$
As was mentioned before Proposition 2.11, in [LeB84, Mas85] the induced CR structure on $\mathcal{T}^{[g]}$ is obtained by identifying it with $\mathcal{N}^{\mathbb{C}}\backslash\mathcal{N}^{\mathbb{R}}.$ The 1-form $\omega^{1}+\mathrm{i}\omega^{2}$ on $\mathcal{T}^{[g]}$ is the horizontal lift of the dual $\mathrm{i}$-eigenspace for the almost complex structure on positively oriented (future time-like) planes using any Weyl connection $\nabla^{[g]}$. The 1-form $\theta^{1}+\mathrm{i}\theta^{2}$ is the canonical holomorphic 1-form on $\mathrm{SO}^{o}(p,q)/\mathrm{SO}(2,\mathbb{R})$ that is invariant under the action of $\mathrm{SO}(2,\mathbb{R}).$ It is straightforward to show that the resulting complex structure on the contact distribution is independent of the choice of $\nabla^{[g]}.$
@@
As was mentioned before, in terms of the adapted coframe (2.17) on $\mathcal{T}^{[g]},$ the value $\lambda=0$ in (2.11), corresponds to the fibers of $\mathcal{T}^{[g]}\to M.$ Thus, by Proposition 2.9 one knows a priori that the pull-back of the binary quartic of $\mathcal{T}^{[g]}$ to $\mathcal{C}$ vanishes at $\lambda=0.$ The fact that $\mathcal{T}^{[g]}$ has a repeated root of multiplicity at least four is an analogue of the fact that the Weyl curvature of 4-dimensional Lorentzian Fefferman conformal structures, defined for any CR 3-manifold, has Petrov type $\mathsf{N}$ [Lew88]. Similarly, it turns out the twistor bundle defined in Remark 2.10 carries a nondegenerate para-CR structure whose fundamental quartic satisfies the same properties as in Proposition 2.11.
@@ 3.1. Projective structures in dimension three
A projective structure []delimited-[][$\nabla$][ ∇ ] on a manifold $M$ is given by an equivalence class of torsion-free connections on $M$ whose geodesics coincide as unparametrized curves. If $M$ is equipped with an orientation, then []delimited-[][$\nabla$][ ∇ ] is called an oriented projective structure. Let $\gamma(t)=(x^{i}(t))\subset M$ be a geodesic for [].delimited-[][$\nabla$].[ ∇ ] . Denoting its tangent vector field as $\dot{\gamma}(t)=(x_{t}^{i}(t)),$ if $\nabla,\tilde{\nabla}\in[\nabla],$ then there is a function $v(t)$ along $\gamma(t)$ such that
(3.1) (~$(\tilde{\nabla}_{\dot{\gamma}(t)}-\nabla_{\dot{\gamma}(t)})\dot{\gamma}(t)=v(t)\dot{\gamma}(t)\Rightarrow(\tilde{\Gamma}^{i}_{jk}-\Gamma^{i}_{jk})x_{t}^{j}(t)x_{t}^{k}(t)=v(t)x_{t}^{i}(t)\Rightarrow\tilde{\Gamma}^{i}_{jk}=\Gamma^{i}_{jk}+\delta^{i}_{j}f_{k}+\delta^{i}_{k}f_{j},$
@@
Solving the equivalence problem of projective structures goes back to Cartan [Car24]. The solution is given as Cartan geometries of type $(\mathrm{SL}(n+1,\mathbb{R}),P_{1}),$ where $P_{1}=\mathrm{GL}(n,\mathbb{R})\ltimes\mathbb{R}^{n}\subset\mathrm{SL}(n+1,\mathbb{R})$ is the parabolic subgroup that stabilizes a line in $\mathbb{R}^{n+1}.$ In the case of oriented projective structure $P_{1}=\mathrm{GL}^{+}(n,\mathbb{R})\ltimes\mathbb{R}^{n}$ is the stabilizer of a ray. In our case, we are interested in 3-dimensional projective structures, which correspond to regular and normal Cartan geometries $(\pi\colon\mathcal{P}\to M,\varphi)$ of type $(\mathrm{SL}(4,\mathbb{R}),P_{1}).$ The Cartan connection and Cartan curvature can be written as
(3.2) $\varphi=\begin{pmatrix}-\eta^{i}_{i}&\mu_{2}&\mu_{1}&\mu_{0}\\ \tau^{2}&\eta^{2}_{2}&\eta^{2}_{1}&\eta^{2}_{0}\\ \tau^{1}&\eta^{1}_{2}&\eta^{1}_{1}&\eta^{1}_{0}\\ \tau^{0}&\eta^{0}_{2}&\eta^{0}_{1}&\eta^{0}_{0}\end{pmatrix},\quad\boldsymbol{\varphi}=\mathrm{d}\varphi+\varphi{\,{\wedge}\;}\varphi=\begin{pmatrix}0&\boldsymbol{\mu}_{2}&\boldsymbol{\mu}_{1}&\boldsymbol{\mu}_{0}\\ 0&\boldsymbol{\eta}^{2}_{2}&\boldsymbol{\eta}^{2}_{1}&\boldsymbol{\eta}^{2}_{0}\\ 0&\boldsymbol{\eta}^{1}_{2}&\boldsymbol{\eta}^{1}_{1}&\boldsymbol{\eta}^{1}_{0}\\ 0&\boldsymbol{\eta}^{0}_{2}&\boldsymbol{\eta}^{0}_{1}&\boldsymbol{\eta}^{0}_{0}\end{pmatrix},$
@@
(3.3) $\boldsymbol{\eta}^{i}_{j}=\textstyle{\frac{1}{2}}W^{i}_{jkl}\tau^{k}{\,{\wedge}\;}\tau^{k},\quad\boldsymbol{\mu}_{j}=\textstyle{\frac{1}{2}}W_{jkl}\tau^{k}{\,{\wedge}\;}\tau^{k},\quad W^{i}_{jik}=0,\quad W^{i}_{[ijk]}=0,\quad W_{[ijk]}=0.$
The structure functions $W^{i}_{jkl}$ on $\mathcal{P}$ are referred to the projective Weyl curvature whose vanishing imply flatness, i.e. the projective structure is locally equivalent to projective lines on $\mathbb{P}^{3}.$
The 1-forms $\tau^{0},\tau^{1},\tau^{2}$ are semi-basic with respect to the fibration $\mathcal{P}\to M.$ The leaves of the integrable Pfaffian system $\{\tau^{0},\tau^{1},\eta^{0}_{2},\eta^{1}_{2}\}$ on $\mathcal{P}$ project to the geodesics of []delimited-[][$\nabla$][ ∇ ] on $M.$
[…]

### v3

@@ 2.4. Twistor CR manifolds and their Petrov type
Define $\mathcal{T}^{[g]}=\mathcal{R}^{\uparrow}T^{*}M,$ where $\mathcal{R}^{\uparrow}T^{*}M$ denotes the bundle of rays in $T^{*}M$ with respect to a Riemannian conformal structure, i.e. the two-fold cover of $\mathbb{P}T^{*}M$, while in the Lorentzian signature it denotes the rays of future time-like 1-forms. Thus, the twistor bundle $\widetilde{\mu}\colon\mathcal{T}^{[g]}\to M$ is a 5-manifold whose fibers are the 2-sphere, $\mathbb{S}^{2},$ if $[g]$ is Riemannian or the 2-disk, $\mathbb{D}^{2},$ if $[g]$ is Lorentzian. To view $\hat{\mu}\colon\mathcal{C}\to\mathcal{T}^{[g]}$ as an associated bundle of $\mu\colon\mathcal{C}\to M,$ recall that $(\mathcal{C}\to M,\psi)$ has type $(\mathrm{SO}^{o}(p+1,q+1),Q_{1}),$ where $(p,q)$ is either (3,0) or (2,1), $Q_{1}=Q_{0}\ltimes\mathbb{R}^{3},$ $Q_{0}=\mathrm{CO}^{o}(p,q)\cong R^{+}\times\mathrm{SO}^{o}(p,q)$ is the structure group, $\mathrm{SO}^{o}(2,1)$ is the identity component of $\mathrm{SO}(2,1)$ and $\mathrm{SO}^{o}(3,\mathbb{R})=\mathrm{SO}(3,\mathbb{R}).$ The stabilizer of a positively oriented (time-like) plane defines a subgroup $S\subset Q_{1}.$ Thus, one obtains that $\hat{\mu}\colon\mathcal{C}\to\mathcal{T}^{[g]}$ is a principal $S$-bundle i.e.
(2.14) $\mathcal{T}^{[g]}\cong\mathcal{C}/\penalty 50S:=\mathcal{C}\times_{Q_{1}}(Q_{1}/\penalty 50S),\quad\text{where}\quad S=\mathrm{CO}(2,\mathbb{R})\ltimes\mathbb{R}^{3}.$
@@
[…]
@@
In Lorentzian signature, the action of $Q_{1}$ on each oriented projectivized cotangent space has five orbits given by future and past rays of time-like 1-forms, rays of space-like 1-forms, and rays of future and past null 1-forms. One can define a twistor bundle using space-like planes in each tangent space equipped with their canonically induced almost para-complex structure. The construction goes through as before with the only difference being the fibers which are $\mathrm{SO}(2,1)/\penalty 50\mathrm{SO}(1,1)\cong\mathbb{R}\times\mathbb{R}\mathbb{P}^{1}.$
As was mentioned above, $\mathcal{T}^{[g]}$ was originally defined as a connected component of the set of complexified null 2-planes without any real null 2-planes, defined in (2.5), i.e. $\mathcal{N}^{\mathbb{C}}$ in the Riemannian signature and a connected component of $\mathcal{N}^{\mathbb{C}}\backslash\mathcal{N}^{\mathbb{R}}$ in the Lorentzian signature. The two definitions are equivalent since there is a bijection between (time-like) planes equipped with their canonically induced almost complex structure and projective classes of complex null 1-forms $\omega_{z}]:=[z_{0}\omega^{0}+z_{1}\omega^{1}+z_{2}\omega^{2}]$ in (2.5). To see this, note that given a positive (time-like) plane $p\in\mathrm{Gr}^{+}(2,T_{x}M)$, one can find a coframe $(\omega^{0},\omega^{1},\omega^{2})$ that satisfies (2.2) and $p=\mathrm{Ker}\{\omega^{0}\}.$ The induced $\mathrm{SO}(2,\mathbb{R})$ action on $(\omega^{1},\omega^{2})^{t}$ defines an almost complex structure whose dual $\mathrm{i}$-eigenspace is $\omega^{1}+\mathrm{i}\omega^{2},$ which is a null complex 1-form. Conversely, given a projective class of a complex null 1-form, $\omega_{z}]:=[z_{0}\omega^{0}+z_{1}\omega^{1}+z_{2}\omega^{2}],$ $z_{0}^{2}+\varepsilon z_{1}^{2}+\varepsilon z_{2}^{2}=0,$ one can find $A^{i}_{j}]\in\mathrm{SO}^{o}(p,q)$ such that $\omega_{z}]=[\tilde{\omega}^{1}+\mathrm{i}\tilde{\omega}^{2}]$ where $\tilde{\omega}^{i}=A^{i}_{j}\omega^{j}.$ More explicitly, examining the case $(p,q)=(3,0)$ over the affine chart $z_{2}=\mathrm{i},$ one has $\omega_{z}]=[z_{0}\omega^{0}+z_{1}\omega^{1}+\mathrm{i}\omega^{2}]$ and $z_{0}^{2}+z_{1}^{2}=1.$ Using the 2-to-1 isogeny $\mathrm{SU}(2)\to\mathrm{SO}(3,\mathbb{R}),$ the coframe transformation $\omega^{i}\to\tilde{\omega}^{i}$ is
@@
where $|a|^{2}+|b|^{2}=1.$ Finding $a,b\in\mathbb{C}$ such that $\tilde{\omega}^{1}+\mathrm{i}\tilde{\omega}^{2}]=[z_{0}\omega^{0}+z_{1}\omega^{1}+\mathrm{i}\omega^{2}]$ results in a set of polynomial relations that can be easily solved for any $z_{0},z_{1}\in\mathbb{C}$ satisfying $z_{0}^{2}+z_{1}^{2}=1.$ This shows that $\omega_{z}$ gives the dual $\mathrm{i}$-eigenspace for the almost complex structure on $\mathrm{Ker}\{\tilde{\omega}^{0}\}.$
@@
To show that $\mathcal{T}^{[g]}$ carries a CR structure, we follow [SY89], which is different from [LeB84] and is based on exploiting the embedding of Lie algebras $\mathfrak{so}(p+1,q+1)\hookrightarrow\mathfrak{su}(2,2),$ where $p+q=3,$ given explicitly in (2.16). Firstly, using the description $\mathcal{T}^{[g]}=\mathcal{C}/\penalty 50S$ in (2.14) together with the Cartan connection (2.1), one obtains that the 1-forms $(\omega^{0},\omega^{1},\omega^{2},\theta^{1},\theta^{2})$ are semi-basic with respect to the fibration $\mathcal{C}\to\mathcal{T}^{[g]}.$ Furthermore, by the structure equations (2.3), one knows
@@
where $\omega=\omega^{1}+\mathrm{i}\omega^{2}$ and $\theta=\theta^{1}+\mathrm{i}\theta^{2}.$ The conformal class of Hermitian inner products $h]\subset\mathrm{Sym}^{2}(\mathscr{C}^{*})$ is given by $h=\tfrac{\mathrm{i}}{2}(\omega\overline{\theta}-\overline{\omega}\theta).$ Using an adapted coframe $(\omega^{0},\omega,\theta,\overline{\omega},\overline{\theta}),$ one can find the Cartan connection for the induced nondegenerate CR structure on $\mathcal{T}^{[g]}.$ Namely, the Lie algebra embedding $\mathfrak{so}(p+1,q+1)\hookrightarrow\mathfrak{su}(2,2)$ results in the injection $\iota\colon S\to P,$ where $P\in\mathrm{SU}(2,2)$ is the parabolic subgroup in Theorem 2.5 and $S$ is given in (2.14). Defining $\mathcal{Q}=\mathcal{C}\times_{S}P,$ one obtains a natural inclusion $\iota\colon\mathcal{C}\to\mathcal{Q}$ between the Cartan geometries $(\mathcal{C}\to M,\psi)$ and $(\mathcal{Q}\to\mathcal{T}^{[g]},\gamma)$ via which
@@
are defined in terms of the entries of the Cartan connection (2.1) for the 3-dimensional conformal structure $[g]$. Using $\iota^{*}\gamma$ and the structure equations (2.3) and (2.8), the coefficients $C_{0},\cdots,C_{4}$ of the holomorphic binary quartic (2.9) on $\mathcal{T}^{[g]},$ pulled-back via $\iota,$ are found to be
@@
The Cotton-York entry $Y_{012}$ is identically zero everywhere on $\mathcal{C}$ if and only if the conformal structure is flat. As a result, the CR structure on $\mathcal{T}^{[g]}$ is flat if and only if the conformal structure $[g]$ is flat. Otherwise, from (2.18), it follows that the binary quartic (2.9) has a repeated root of multiplicity four. Using the parametrization of $\alpha$-planes (2.11) and the quartic (2.12), it follows that the repeated root corresponds to the $\alpha$-plane for which $a_{1}:a_{2}]=[0:1],$ i.e. the contact planes $\mathrm{Ker}\{\omega^{0},\omega^{1},\omega^{2}\},$ whose integral surfaces are the fibers of $\mathcal{T}^{[g]}\to M.$
As was mentioned before Proposition 2.11, in [LeB84, Mas85] the induced CR structure on $\mathcal{T}^{[g]}$ is obtained by identifying it with $\mathcal{N}^{\mathbb{C}}\backslash\mathcal{N}^{\mathbb{R}}.$ The 1-form $\omega^{1}+\mathrm{i}\omega^{2}$ on $\mathcal{T}^{[g]}$ is the horizontal lift of the dual $\mathrm{i}$-eigenspace for the almost complex structure on positively oriented (future time-like) planes using any Weyl connection $\nabla^{[g]}$. The 1-form $\theta^{1}+\mathrm{i}\theta^{2}$ is the canonical holomorphic 1-form on $\mathrm{SO}^{o}(p,q)/\penalty 50\mathrm{SO}(2,\mathbb{R})$ that is invariant under the action of $\mathrm{SO}(2,\mathbb{R}).$ It is straightforward to show that the resulting complex structure on the contact distribution is independent of the choice of $\nabla^{[g]}.$
@@
As was mentioned before, in terms of the adapted coframe (2.17) on $\mathcal{T}^{[g]},$ the value $\lambda:=a_{1}/a_{2}=0$ in the parametrization (2.11), corresponds to the fibers of $\mathcal{T}^{[g]}\to M.$ Thus, by Proposition 2.9 one knows a priori that the pull-back of the binary quartic of $\mathcal{T}^{[g]}$ to $\mathcal{C}$ vanishes at $\lambda=0.$ The fact that $\mathcal{T}^{[g]}$ has a repeated root of multiplicity at least four is an analogue of the fact that the Weyl curvature of 4-dimensional Lorentzian Fefferman conformal structures, defined for any CR 3-manifold, has Petrov type $\mathsf{N}$ [Lew88]. Similarly, it turns out the twistor bundle defined in Remark 2.10 carries a nondegenerate para-CR structure whose fundamental binary quartic satisfies the same properties as in Proposition 2.11.
@@ 3.1. Projective structures in dimension three
A projective structure $\nabla$] on a manifold $M$ is given by an equivalence class of torsion-free connections on $M$ whose geodesics coincide as unparametrized curves. If $M$ is equipped with an orientation, then $\nabla$] is called an oriented projective structure. Let $\gamma(t)=(x^{i}(t))\subset M$ be a geodesic for $\nabla$]. Denoting its tangent vector field as $\dot{\gamma}(t)=(x_{t}^{i}(t)),$ if $\nabla,\tilde{\nabla}\in$$\nabla$], then there is a function $v(t)$ along $\gamma(t)$ such that
(3.1) $(\tilde{\nabla}_{\dot{\gamma}(t)}-\nabla_{\dot{\gamma}(t)})\dot{\gamma}(t)=v(t)\dot{\gamma}(t)\Rightarrow(\tilde{\Gamma}^{i}_{jk}-\Gamma^{i}_{jk})x_{t}^{j}(t)x_{t}^{k}(t)=v(t)x_{t}^{i}(t)\Rightarrow\tilde{\Gamma}^{i}_{jk}=\Gamma^{i}_{jk}+\delta^{i}_{j}f_{k}+\delta^{i}_{k}f_{j},$
@@
Solving the equivalence problem of projective structures goes back to Cartan [Car24]. The solution is given as Cartan geometries of type $(\mathrm{SL}(n+1,\mathbb{R}),P_{1}),$ where $P_{1}=\mathrm{GL}(n,\mathbb{R})\ltimes\mathbb{R}^{n}\subset\mathrm{SL}(n+1,\mathbb{R})$ is the parabolic subgroup that stabilizes a line in $\mathbb{R}^{n+1}.$ In the case of oriented projective structures $P_{1}=\mathrm{GL}^{+}(n,\mathbb{R})\ltimes\mathbb{R}^{n}$ is the stabilizer of a ray. We are interested in 3-dimensional projective structures, which correspond to regular and normal Cartan geometries $(\pi\colon\mathcal{P}\to M,\varphi)$ of type $(\mathrm{SL}(4,\mathbb{R}),P_{1}).$ The Cartan connection and Cartan curvature, which can be found in [KN64], are given by
(3.2) $\varphi=\begin{pmatrix}-\eta^{i}_{i}&\mu_{2}&\mu_{1}&\mu_{0}\\ \tau^{2}&\eta^{2}_{2}&\eta^{2}_{1}&\eta^{2}_{0}\\ \tau^{1}&\eta^{1}_{2}&\eta^{1}_{1}&\eta^{1}_{0}\\ \tau^{0}&\eta^{0}_{2}&\eta^{0}_{1}&\eta^{0}_{0}\end{pmatrix},\quad\Psi=\mathrm{d}\varphi+\varphi{\,{\wedge}\;}\varphi=\begin{pmatrix}0&\Gamma_{2}&\Gamma_{1}&\Gamma_{0}\\ 0&\Lambda^{2}_{2}&\Lambda^{2}_{1}&\Lambda^{2}_{0}\\ 0&\Lambda^{1}_{2}&\Lambda^{1}_{1}&\Lambda^{1}_{0}\\ 0&\Lambda^{0}_{2}&\Lambda^{0}_{1}&\Lambda^{0}_{0}\end{pmatrix}$,
@@
(3.3) $\Lambda^{i}_{j}=\textstyle{\frac{1}{2}}W^{i}_{jkl}\tau^{k}{\,{\wedge}\;}\tau^{l},\quad\Gamma_{j}=\textstyle{\frac{1}{2}}W_{jkl}\tau^{k}{\,{\wedge}\;}\tau^{l},\quad W^{i}_{jik}=0,\quad W^{i}_{[jkl]}=0,\quad W_{[ijk]}=0.$
The structure functions $W^{i}_{jkl}$ on $\mathcal{P}$ are the entries of the projective Weyl curvature whose vanishing imply flatness, i.e. the projective structure is locally equivalent to the one defined by projective lines in $\mathbb{P}^{3}.$
The 1-forms $\tau^{0},\tau^{1},\tau^{2}$ are semi-basic with respect to the fibration $\mathcal{P}\to M.$ Furthermore, the leaves of the integrable Pfaffian system $\mathcal{H}=\{\tau^{0},\tau^{1},\eta^{0}_{2},\eta^{1}_{2}\}$ on $\mathcal{P}$ project to the geodesics of $\nabla$] on $M,$ which in the model space $\mathbb{P}^{3}$ can be viewed as follows: if $\ell\subset\mathbb{P}^{3}$ is the the projectivization of the 2-plane $\langle e_{3},e_{2}\rangle\subset\mathbb{R}^{4},$ where $(e_{3},e_{2},e_{1},e_{0})$ is the standard basis for $\mathbb{R}^{4},$ the Lie algebra of the stabilizer of $\ell$ is given by $\mathfrak{sl}(4,\mathbb{R})$-valued matrices for which the bottom left $2\times 2$ block is zero, i.e. the stabilizer is the parabolic subalgebra $\mathfrak{p}_{2}\subset\mathfrak{sl}(4,\mathbb{R})$ defined by crossing the second node in the Dynkin diagram of $A_{3}.$ In terms of the Cartan connection (3.2), the stabilizer of $\ell$ corresponds to the Pfaffian system $\mathcal{H}.$ In the curved case, this can be seen as follows: the 5-dimensional leaf space of the Pfaffian system $\mathcal{K}=\{\tau^{0},\tau^{1},\tau^{2},\eta^{0}_{2},\eta^{1}_{2}\}$ can be identified with $\tilde{\pi}\colon\mathbb{P}TM\to M,$ since the rank 2 vertical distribution $\mathscr{U}=\mathrm{Ker}\{\tau^{0},\tau^{1},\tau^{2}\}$ together with the tautological rank 3 distribution $\mathscr{T}=\mathrm{Ker}\{\tau^{0},\tau^{1}\}$ satisfy $\mathscr{U},\mathscr{T}]=T\mathbb{P}TM.$ Recall that the rank 3 tautological distribution at $p=(x,[v])\in\mathbb{P}TM,$ where $x\in M$ and $v]\subset T_{x}M$ is a line, is defined as $\mathscr{T}=\tilde{\pi}_{*}^{-1}([v])\subset T_{p}\mathbb{P}TM.$ Thus, the distributions $(\mathscr{U},\mathscr{T})$ define the canonical multi-contact structure on $\mathbb{P}TM,$ i.e. the leaf space of $\mathcal{K}$ is locally equivalent to $\mathbb{P}TM.$ If $\tilde{\mathscr{E}}\subset TTM$ denotes the rank 3 horizontal distribution defined by $\nabla\subset[\nabla],$ then $\ell:=\mathscr{T}\cap\tilde{\pi}_{*}(\tilde{\mathscr{E}})\subset T\mathbb{P}TM$ is the line field $\ell=\mathrm{Ker}\mathcal{H}=\mathrm{Ker}\{\tau^{0},\tau^{1},\eta^{0}_{2},\eta^{1}_{2}\},$ referred to as the geodesic spray of $\nabla,$ which is invariant of the choice of connection $\nabla,$ and, via projection to $M,$ its integral curves exhaust the geodesics of $\nabla$] in a one-to-one fashion.
The geodesic equation for any representative connection $\nabla\in[\nabla]$ gives a triple of second order ODEs $x^{i}_{tt}=-\Gamma^{i}_{jk}x^{j}_{t}x^{k}_{t}$, where the independent parameter $t$ is defined up to homotheties and translations. Considering geodesics as unparametrized curves, a local coordinate on $M$ can be taken as the independent parameter, which allows one to arrive at a pair of second order ODEs, defined up to point transformations, as follows. Working locally with an open set of geodesics along which $x^{0}_{t}\neq 0,$ one can set $t=t(x_{0})$ and $x^{i}=x^{i}(x^{0})$ for $i=1,2,$ along such geodesics. Defining $\dot{x}^{i}=\tfrac{\mathrm{d}x^{i}}{\mathrm{d}x^{0}}=\tfrac{x^{i}_{t}}{x^{0}_{t}}$ and $\ddot{x}^{i}=\tfrac{\mathrm{d}^{2}x^{i}}{(\mathrm{d}x^{0})^{2}}=\tfrac{1}{x^{0}_{t}}\tfrac{\mathrm{d}}{\mathrm{d}t}(\tfrac{x^{i}_{t}}{x^{0}_{t}})=\tfrac{x^{i}_{tt}}{(x^{0}_{t})^{2}}-\tfrac{x^{i}_{t}x^{0}_{tt}}{(x^{0}_{t})^{3}},$ the geodesic equation reduces to the pair of 2nd order ODEs